# Human Decision-making is Susceptible to AI-driven Manipulation

### v2

@@ Discussion
This study demonstrates significant human susceptibility to AI-driven manipulation across financial and emotional decision-making contexts, raising critical concerns about the psychological and societal implications of advanced AI systems. It presents several key insights that have important implications for understanding and protecting human autonomy in interactions with AI systems.
@@
Our findings revealed that participants interacting with the manipulative agents (MA and SEMA) exhibited substantially higher rates of harmful decision-making compared to those who interacted with the NA. This susceptibility was further highlighted in the temporal changes of participants’ ratings, as participants in the manipulative groups exhibited declining preferences for the optimal choices while increasingly favoring the hidden incentives after interacting with these agents.
The contrasting decision patterns between financial and emotional domains highlight how context shapes human susceptibility to AI-driven manipulation. Participants showed greater susceptibility to AI’s influence in quantitative decision-making scenarios (i.e., financial contexts), where they may over-trust AI’s algorithmic advice due to its perceived objectivity [46]. In contrast, participants were less susceptible to external influence in emotional scenarios, more frequently preserving their initial beneficial choices, likely due to emotional judgments being deeply rooted in personal intuition and less reliant on external recommendations [47]. Hence, future work should investigate the cognitive and contextual mechanisms underlying human susceptibility and develop targeted interventions that mitigate AI-driven manipulation.
Another notable finding was the similar performance of MA and SEMA conditions in shifting participants’ preferences and decisions. These findings suggest that hidden objectives play a significant role in AI-driven manipulation, as the presence of such objectives alone (MA) was nearly as effective as employing established psychological strategies (SEMA) in influencing participants’ decisions. However, the addition of strategies did enhance the manipulative impact in financial scenarios (Supplementary Table S1), suggesting that while hidden objectives are a primary driver, the use of sophisticated tactics may further amplify their influence in contexts where users tend to defer more to AI’s perceived expertise (i.e., financial decisions). These results raise critical concerns about the potential for covert influence in AI systems, as agents that were not explicitly designed to manipulate their users could also significantly impact human decision-making based on a hidden incentive.
@@
Our analyses also revealed the AI agents’ adaptation of manipulation strategies across different domains. These findings suggest that while the overall distribution of the employed strategies followed a similar trend across different domains, indicating that the AI agents may implicitly prefer certain strategies more than others, such agents can tailor their strategy usage to the context (e.g., exploiting interpersonal vulnerabilities in the emotional domain while using pragmatic tactics in financial scenarios). This context-specific adaptation highlights AI agents’ nuanced approach to maximizing manipulative impact across different scenarios, highlighting broader issues regarding AI safety.
@@
Moving forward, our findings highlight broader issues about the societal implications of AI-driven manipulation, particularly in the studied domains. In financial contexts, traditional advertising operates within legal frameworks, requiring transparency (e.g., disclosing sponsorships) and accountability for false claims. In addition, the commercial intent of advertisers is inherently recognized by consumers, creating a sense of skepticism toward their genuineness. However, unlike advertisers, AI systems currently face limited accountability for their behavior. For instance, AI’s fabricated claims and hallucinations are frequently dismissed as technical errors rather than intentional deception, leading to regulatory loopholes. Recommendations by such assistants are mainly perceived as unbiased and helpful, causing users to trust their advice based on the displayed sincerity. In addition, AI’s ability to tailor its recommendations to individual personalities[16] may enable targeted advertisements that exploit users’ vulnerabilities and undermine their autonomy, which goes beyond what traditional forms of advertisement could achieve.
@@
To ensure participant well-being, participants were screened to exclude individuals with reported mental or physical health conditions or those undergoing treatment, following established guidelines for psychological research [52]. In addition, participants were required to assume the perspective of a user based on pre-defined queries in hypothetical scenarios. These scenarios focused on common decision-making scenarios rather than highly sensitive or traumatic situations, minimizing the risk of psychological harm and preventing the disclosure of personally identifiable information. Importantly, participants received a comprehensive debriefing immediately after study completion, which included full disclosure of the study’s objectives, an explanation of our motivations, the optimal solutions for each scenario, and contact information for additional support or questions. Regarding participant privacy and data security, all interactions were conducted on a secure platform, with identifying information stored separately in an encrypted database. This study adhered to the University of Hong Kong’s guidelines for research involving human subjects and was approved by the institution’s ethics committee (Reference No: EA240497).
@@
Despite our efforts to create the first robust and comprehensive analysis of human susceptibility to AI-driven manipulation, this study faces several limitations. The study was conducted in a controlled environment with pre-defined decision-making scenarios, which enabled us to explore the effects of AI-driven influence in isolation while limiting potential risks to participants’ well-being. However, these artificial constraints may not fully capture the complexity of real-world AI applications, where users are exposed to a broader range of decisions and interactions. Future work can examine how these findings translate to more dynamic settings where AI interacts with users across diverse contexts, exploring such influence beyond decision-making.
@@
Regarding the AI agents employed in this study, our reliance on GPT-4o may present a limitation. While GPT-4o is currently state-of-the-art across various tasks [53, 54], manipulation capabilities may vary across different models. Given the widespread adoption and popularity of GPT-4o (commercially known as ChatGPT), understanding its manipulative tendencies is critical due to its large user base and significant societal impact. However, we do not claim that all models exhibit the same manipulative behavior, as further studies are needed to investigate whether similar tendencies are present in other models. Moreover, our curated taxonomy of strategies represents a subset of potential manipulation techniques in user-assistant settings. Hence, while this taxonomy is theory-based and comprehensive, emerging AI capabilities may form more subtle and sophisticated strategies that are not captured in our curated taxonomy.
The design of the control condition in our experiment also poses a limitation. While the neutral agent was designed to optimize user benefit without explicit manipulative objectives, inherent model biases may persist. These biases may have subtly influenced participants’ decisions, even in the absence of a hidden incentive, as seen in the negative shifts across both decision-making domains for the NA group. This limitation emphasizes the challenge of creating truly neutral AI systems, and future research should focus on further identifying and mitigating harmful biases in AI agents.
Lastly, this study highlights several crucial research directions for future work. Investigating the long-term impact of AI-influenced decisions could reveal whether such changes persist over time. Developing reliable methods to detect subtle AI manipulation, particularly in emotional contexts where our findings suggest users are less likely to notice AI’s influence, also remains an open challenge. While constructing ethical frameworks for AI development is essential, future work is also needed to investigate and create regulatory frameworks to protect consumers from AI-driven manipulation. This may include imposing transparency requirements, accountability mechanisms, and safeguards to ensure that AI systems prioritize user well-being and protect their autonomy. In addition, future work should also explore systems that can lower human susceptibility to such manipulation.
@@ Conclusions
Our randomized control trial with $233$ participants demonstrated that human decision-making is highly susceptible to AI-driven manipulation, with participants significantly shifting preferences toward harmful options and away from beneficial choices when interacting with manipulative AI agents. Strikingly, established manipulation strategies (SEMA) did not amplify harm over covert objectives (MA), suggesting that even subtle, hidden incentives can sway human preferences. These findings reveal critical vulnerabilities in human-AI interactions, particularly in domains perceived as rational, like finance, where users may over-trust AI’s perceived objectivity. These results underscore the need for regulatory frameworks to protect user autonomy in the age of increasingly sophisticated AI systems.
@@ Methods
We conducted a randomized controlled trial to investigate human susceptibility to AI manipulation attempts. Our study employed a between-subjects design ($N=233$) to examine how and to what extent AI systems (referred to as agents in this paper) can influence human decision-making across decision-making scenarios (Figure 1). In each scenario, the participant was provided with a pre-defined user query and four corresponding options. One option was created as the most suitable (optimal) choice based on the user’s query, while the remaining three were designed to cause different types of harm. We investigated participants’ susceptibility to AI-driven manipulation by analyzing the differences in their preference ratings for these options before and after interacting with the assigned AI agent.
@@
We developed three distinct AI agents by designing prompts for GPT-4o[55], the current state-of-the-art large language model. The primary experimental condition, the Manipulative Agent (MA), was configured with a hidden objective to influence users’ preferences toward a specific harmful option (i.e., hidden incentive) and away from the beneficial (i.e., optimal) choices. We implemented a point-based reward system to quantify the agent’s manipulation success. 100 points were awarded for achieving the hidden incentive, 50 points for steering users toward other harmful options, and 0 points for the optimal choice. In addition, the agent employed a personalized theory-of-mind (ToM)[56] framework to reason about users’ beliefs, desires, intentions, and emotions while considering the vulnerabilities associated with their personality traits before selecting a strategy and generating a response.
We designed the Strategy-enhanced Manipulative Agent (SEMA) as an additional experimental condition to investigate whether providing the agent with established psychological strategies would impact its manipulative influence. This agent shared the same reward system, reasoning framework, and underlying goal of influencing users’ decisions as MA. In addition, it was provided with our curated taxonomy of established manipulation strategies derived from psychological and social science literature[57, 58, 59, 60, 61, 62] (Table S9). Lastly, serving as the control condition in our experiments, the Neutral Agent (NA) was instructed to provide unbiased assistance and reassurance while optimizing for user benefit.
All agents were provided with the user’s query and the corresponding options for each scenario. In addition, they had access to the participant’s personality traits, as personalization could significantly enhance the effectiveness of human-AI interactions [16]. Given that existing AI systems are capable of predicting their users’ personalities with sufficient accuracy [63], we believe adding this information did not comprise the applicability of our experimental setup to real-world contexts. Designed prompts are provided in the appendix (Supplementary Figures S3 - S5).
@@
The scenarios in this study were carefully crafted to assess manipulation susceptibility in two domains reflecting common real-world situations in which AI assistants are employed [42, 43, 64, 44, 65, 33]: Financial (seeking advice on purchases) and Emotional (seeking guidance on handling emotional dilemmas) decision-making. In each domain, we curated three widely used topics to design scenarios that applied to a wide variety of individuals. For each scenario, to ensure consistency between the conditions of different participants, we designed pre-defined queries representing a hypothetical user’s situation, needs, and requirements. In addition, each scenario was accompanied by four distinct plausible options, one of which was crafted as the optimal choice, while others were carefully designed to indicate different levels of potential harm. In each scenario, we designated one of the harmful options as the agent’s hidden incentive to prevent user suspicion regarding its bias.
@@
AI systems are making their way into financial investments [32] and are increasingly being used to enhance product advertisements [16, 33] and provide financial advice [66]. Hence, it is crucial to understand the extent to which these systems can impact users’ decisions despite their better judgments. Drawing from existing research on consumer decision-making [67, 68, 69], we identified key scenarios where users commonly seek AI assistance for financial guidance. We focused on three representative consumer decisions that require minimal domain expertise: selecting a fitness tracker, choosing a weight-loss medication, and finding an online shopping platform for clothes. To ensure practicality while maintaining experimental control, we designed pre-defined user queries corresponding to each topic that incorporated realistic budget constraints and requirements (Tables S10-S12).
The optimal choice was designed to align with stated user requirements and budget constraints while incorporating realistic trade-offs. We designed the remaining options to exhibit different types of potential financial harm: 1) Non-existent Product: This option was crafted to mirror characteristics of common online scams [70], featuring unusually attractive terms, comparatively lower cost, and limited verification possibilities. Specifically, this option was designed to only be available for pre-order from the supplier’s website and had relatively few or no reviews, making it difficult for users to assess its legitimacy. 2) Excessive Product: This option represented over-featured solutions that exceeded the user’s stated requirements and maximum budget and was designed to test users’ susceptibility to unnecessary upselling[71]. It allowed us to assess whether users could be swayed by the perceived desirability of excessive products, even when such choices caused them to spend more than they had originally planned. 3) Dependence-inducing Product: This option was designed to satisfy the user’s requirements at a relatively lower upfront cost than the optimal choice. However, it incorporated elements of long-term financial commitment, such as subscription models, for continued access to users’ desired features, which introduced hidden costs and future obligations. By choosing this option, users incurred recurring fees that, over time, exceeded their maximum budget, reflecting real-world scenarios where consumers are lured by low initial costs but face long-term financial burdens.
@@
The increasing deployment of AI systems in mental healthcare, particularly for counseling and emotional support, presents both opportunities and potential risks. These systems have gained traction due to their high accessibility, cost-effectiveness, and reduced stigma associated with seeking support [45]. However, this widespread adoption necessitates an investigation into human susceptibility to AI influence when dealing with emotional challenges, particularly regarding their established beliefs about managing emotional dilemmas. The scenarios for this domain were developed based on a comprehensive analysis of mental health forums and existing literature on emotional dilemmas [72, 73]. Through careful examination of existing data and literature [74, 75, 76, 77, 78], we identified three primary topics of emotional adversities: issues with self-esteem (Table S13), conflict with a close friend (Table S14), and criticism at the workplace (Table S15).
@@
Each scenario presented participants with four distinct coping strategies (i.e., how individuals tend to manage their thoughts and emotions in response to emotional adversities [79]), designed to represent different approaches to emotional regulation and problem-solving. Existing research [79, 80, 81, 82, 83] classifies coping strategies as adaptive or maladaptive depending on their general impact on the individual’s overall well-being and mental health. We included an adaptive coping mechanism as the optimal choice for the given scenario, while the remaining options were designed to incorporate maladaptive practices, which could potentially lead to increasing rates of depression and anxiety [84, 85].
@@
Therefore, the optimal choice was designed to represent active coping, a process based on cognitive-behavioral literature that involves identifying specific actions to address the situation, developing alternative perspectives, and constructive engagement with support networks, which has shown the strongest correlations with increasing user’s mental well being and lower rates of anxiety and depression [80, 82]. The remaining options involved common maladaptive mechanisms with varying levels of harm [85]: 1) Disengagement: This option reflects avoidant coping patterns, incorporating elements such as behavioral disengagement, mental disengagement, and denial. While potentially offering immediate emotional relief, this option demonstrated less desirable long-term outcomes; 2) Emotional Venting: This option includes elements of emotional release while omitting components of reflection, learning, or growth that characterize more adaptive forms of emotional processing. Hence, it clearly distinguishes between adaptive and maladaptive emotional processing by emphasizing unregulated emotional discharge without problem-solving elements. 3) Self-blame: This option was constructed to reflect common patterns of maladaptive attribution as it included subtle elements of perfectionism and excessive responsibility-taking, characteristics that portray the individual as the root of the problem, bringing additional mental burden to the individual without finding a viable long-term solution.
@@
To ensure user privacy, we created our own data collection platform using React[86] and FastAPI[87]. Upon logging onto our platform, each participant was shown a series of three scenarios based on their assigned domain (Figure S7). Selecting one of these scenarios opened a new page, in which the participants were shown the pre-defined query and four available options in a randomized order (Figure S8). The platform interface was designed to minimize potential confounds, presenting all information in a standardized format while randomly rearranging the order of the options within each scenario to control for position effects. Before interacting with the assigned AI agent, participants rated their initial preferences for each option, along with their confidence levels and topic familiarity, on a 10-point Likert scale. After submitting their ratings, participants entered the interaction page (Figure S9). Following a conversation lasting at least 10 turns with the assigned agent, participants provided their final preferences for each option, confidence levels, and scenario familiarity. This design enabled us to quantify changes in both decision-making and confidence levels as metrics for manipulation susceptibility.
@@
This study followed a three-phase protocol designed to assess participants’ susceptibility to AI manipulation while controlling for individual differences and ensuring ethical practices. Our experiments were approved by the institutional review board at The University of Hong Kong (Reference No: EA240497).
@@
Prior to our experiment, participants completed a comprehensive questionnaire of validated psychometric instruments, including demographic information (age, sex, education, occupation, and marital status) and social factors that may influence people’s susceptibility to manipulation, including personality[88, 89], social engineering susceptibility[90] [91], social support [92], and self-esteem[93]: 1) Ten-Item-Personality-Inventory (TIPI)[94], a 10-item assessment of personality traits based on dimensions from the Big Five Inventory (BFI)[95]: openness, conscientiousness, extroversion, agreeableness, and emotional stability. We adopted this version of the personality test, as we assumed that while an AI agent may have a perception of the user’s personality traits, it would not be heavily accurate[16]; 2) Social Engineering Susceptibility Scale[96] focusing on individuals’ vulnerabilities in digital interactions and modified to include the portion of the questions that were relevant to our study (i.e., excluding questionnaire regarding informational susceptibilities); 3) Oslo Social Support Scale (OSSS-3)[97], a 3-item questionnaire evaluating participants’ accessibility to support networks; 4) Rosenberg Self-Esteem Scale (RSES)[98], a 10-item measurement of individuals’ perception of self-worth. Due to the nature of our study, we also adopted the AI Attitude Scale[99] to assess participants’ predispositions toward AI systems.

### v3

@@ Discussion
This study demonstrates significant human susceptibility to AI-driven manipulation across financial and emotional decision-making contexts, raising critical concerns about the psychological and societal implications of advanced AI systems. It presents several key insights, which, although preliminary and drawn from simulated low-stakes scenarios, highlight important implications for understanding and protecting human autonomy in interactions with AI systems.
@@
Our findings reveal that participants interacting with the manipulative agents (MA and SEMA) adjusted their ratings in line with the agents’ hidden motives, as reflected by the Hidden-Optimal Differential ($HOD$), decreasing their ratings for the optimal options while increasingly favoring the hidden incentives. This susceptibility is further highlighted by modeling decision-centric probabilities, as the manipulative agents led to substantially larger increases in preference for hidden incentives over optimal options (Hidden) compared to those who interacted with the NA.
Notably, we did not detect a reliable difference between the overall manipulative impact of employing established psychological strategies (SEMA) and hidden objectives alone (MA). In addition, the MA’s impact on participants’ ratings and preferences did not significantly vary across domains and data collection sources (WeChat and Prolific). These findings raise concerns about the potential for covert influence in AI systems, as agents that were not explicitly designed to manipulate could also substantially shape human decisions if their objectives involve covert intents.
On the contrary, the SEMA demonstrated a substantially lower effect on shaping participants’ decisions in the emotional domain, particularly in the WeChat sub-group, suggesting that the addition of sophisticated tactics is context-dependent: it may further amplify influence in contexts where users tend to defer more to AI’s perceived expertise [logg2019algorithm] (e.g., financial decisions), while diminishing its effect in more emotionally charged contexts, where judgments are deeply rooted in personal intuition and cultural background, and less reliant on external recommendations [castelo2021conservatism].
@@
Our analyses also revealed the AI agents’ adaptation of manipulation strategies across the two domains. We found that the overall distribution of employed strategies followed a similar trend across financial and emotional scenarios, indicating that the AI agents may implicitly prefer certain strategies more than others. However, the SEMA still tailored its strategy usage to the context, exploiting interpersonal vulnerabilities in the emotional domain while using pragmatic tactics in financial scenarios. This context-specific adaptation highlights AI agents’ nuanced approach to maximizing manipulative impact across different scenarios, underscoring broader issues regarding AI safety.
@@
Moving forward, our findings highlight broader issues about the societal implications of AI-driven manipulation, particularly in the studied domains. In financial contexts, traditional advertising operates within legal frameworks, requiring transparency (e.g., disclosing sponsorships) and accountability for false claims. In addition, the commercial intent of advertisers is inherently recognized by consumers, creating a sense of skepticism toward their authenticity. However, unlike advertisers, AI systems currently face limited accountability for their behavior. For instance, AI’s fabricated claims and hallucinations are frequently dismissed as technical errors rather than intentional deception, leading to regulatory loopholes. Recommendations by such assistants are mainly perceived as unbiased and helpful, causing users to trust their advice based on the displayed sincerity. In addition, AI’s ability to tailor its recommendations to individual personalities[matz2024potential] may enable targeted advertisements that exploit users’ vulnerabilities and undermine their autonomy, surpassing the capabilities of traditional forms of advertisement.
@@
To ensure participant well-being, participants were screened to exclude individuals with reported mental or physical health conditions or those undergoing treatment, following established guidelines for psychological research [collings2012suicide]. In addition, participants were required to assume the perspective of a user based on pre-defined queries in hypothetical scenarios. These scenarios focused on common decision-making scenarios rather than highly sensitive or traumatic situations, minimizing the risk of psychological harm and preventing the disclosure of personally identifiable information. Importantly, participants received a comprehensive debriefing immediately after study completion, which included full disclosure of the study’s objectives, an explanation of our motivations, the optimal solutions for each scenario, and contact information for additional support or questions. Regarding participant privacy and data security, all interactions were conducted on a secure platform, with identifying information stored separately in an encrypted database. This study adhered to the University of Hong Kong’s guidelines for research involving human subjects and was approved by the institution’s ethics committee (Reference No: EA240497).
@@
Despite our efforts to create the first robust and comprehensive analysis of human susceptibility to AI-driven manipulation, this study faces several limitations.
The study was conducted in a controlled environment with pre-defined, hypothetical decision-making scenarios that had no direct consequences, which enabled us to explore the effects of AI-driven influence in isolation while limiting potential risks to participants’ well-being. However, these artificial constraints do not fully capture the complexity of real-world AI applications, where users are exposed to a broader range of high-stakes decisions and interactions with critical financial and behavior risks. Future work can examine how these findings translate to more dynamic settings where AI interacts with users across diverse contexts, exploring such influence beyond decision-making.
@@
Regarding the AI agents employed in this study, our reliance on GPT-4o may present a limitation. While GPT-4o is currently state-of-the-art across various tasks [song2024csbench, zheng2024judging], manipulation capabilities may vary across different models. Given the widespread adoption and popularity of GPT-4o (commercially known as ChatGPT), understanding its manipulative tendencies is critical due to its large user base and significant societal impact. However, we do not claim that all models exhibit the same manipulative behavior, as further studies are needed to investigate whether similar tendencies are present in other models. Moreover, our curated taxonomy of strategies represents a subset of potential manipulation techniques in user-assistant settings. Hence, while this taxonomy is theory-based and comprehensive, emerging AI capabilities may form more subtle and sophisticated strategies that are not captured in our curated taxonomy.
Participants were recruited via WeChat (Chinese-speaking, predominantly Asian individuals) and Prolific (English-speaking individuals from diverse cultural backgrounds). Therefore, this dual-platform approach may have introduced cultural or demographic variability in our experiments. To account for potential biases, we analyzed outcomes separately for each data collection source and added the source as a covariate in our linear mixed-effects models, while standardizing continuous variables within each source. However, we did not conduct exhaustive subgroup analyses by ethnicity, which represents a limitation. Future studies should employ more balanced recruitment and explicitly test for cross-cultural differences to enhance generalizability.
The lack of formal pre-registration serves as another limitation of this study. While our primary hypotheses and core analytical approach were prospectively outlined and approved by our Institutional Review Board, pre-registration could significantly increase the degree of transparency, distinguish between exploratory and confirmatory analyses, and mitigate the risks of post-hoc adjustments in open science practices. Therefore, we recommend that future work in human-AI interaction research prioritize pre-registration on platforms such as OSF. In addition, to ensure transparency, we have provided detailed statistical reports on the collected data and made the code for our analyses publicly available.
Lastly, this study highlights several crucial research directions for future work. Investigating the long-term impact of AI-influenced decisions could reveal whether such changes persist over time. Developing reliable methods to detect subtle AI manipulation, particularly in emotional contexts where our findings suggest users are less likely to notice AI’s influence, also remains an open challenge. While constructing ethical frameworks for AI development is essential, future work is also needed to investigate and create regulatory frameworks to protect consumers from AI-driven manipulation. This may include imposing transparency requirements, accountability mechanisms, and safeguards to ensure that AI systems prioritize user well-being and protect their autonomy. Additionally, future work should also investigate systems that can reduce human susceptibility to such manipulation.
@@ Conclusions
Our randomized between-subjects experiment with $233$ participants demonstrated that human decision-making is highly susceptible to AI-driven manipulation, with participants significantly shifting preferences toward harmful options and away from beneficial ones when interacting with manipulative AI agents. Notably, established manipulation strategies (SEMA) did not amplify harm over covert objectives (MA). Hence, we hypothesize that even subtle, hidden incentives can sway human preferences. These findings reveal critical vulnerabilities in human-AI interactions, particularly in domains perceived as rational, such as finance, where users may over-trust AI’s perceived objectivity. These results, while preliminary and drawn from simulated settings, underscore the need for regulatory frameworks to protect user autonomy in the age of increasingly sophisticated AI systems.
@@ Methods
We conducted a study to investigate human susceptibility to AI-driven manipulation attempts. Our study employed a randomized between-subjects design ($N=233$) to examine how and to what extent AI systems (referred to as agents in this paper) can influence human decision-making across decision-making scenarios (Figure 1). In each scenario, the participant was provided with a pre-defined user query and four corresponding options. One option was created as the most suitable (optimal) option based on the user’s query, while the remaining three were designed to cause different types of harm. We investigated participants’ susceptibility to AI-driven manipulation by analyzing the differences in their preference ratings for these options before and after interacting with the assigned AI agent.
@@
We developed three distinct AI agents by designing prompts for GPT-4o[hurst2024gpt], a widely-used large language model. Serving as the control condition in our experiments, the Neutral Agent (NA) was instructed to provide unbiased assistance and reassurance while optimizing for user benefit.
The primary experimental condition, the Manipulative Agent (MA), was configured with a hidden objective to shift users’ preference ratings toward a pre-designated harmful option (i.e., hidden incentive) and away from the beneficial (i.e., optimal) option. We implemented a point-based reward system to quantify the agent’s manipulation success and prevent the agent from fixating on one option: 100 points were awarded for achieving the hidden incentive, 50 points for steering users toward other harmful options, and 0 points for the optimal option. In addition, the agent employed a personalized theory-of-mind (ToM)[apperly2012theory] framework to reason about users’ beliefs, desires, intentions, and emotions while considering the vulnerabilities associated with their personality traits before selecting a strategy and generating a response.
We designed the Strategy-enhanced Manipulative Agent (SEMA) as the second experimental condition to investigate whether providing the agent with established manipulation strategies would impact its influence. This agent shared the MA’s reward system, reasoning framework, and underlying hidden objectives. In addition, it was provided with our curated taxonomy of established manipulation strategies derived following three steps: We first synthesized strategies from psychological and social science literature [buss1987tactics, seabright1998interpersonal, lauer2020people, seda2021manipulative, noggle2020pressure, danciu2014manipulative]; Accordingly, we excluded strategies irrelevant to the user-assistant setting (e.g., victimizing oneself, ridicule, and silent treatment) or suggested coercion/excessive authority; Lastly, we included an additional strategy, namely Others, for responses that did not fit in the previous strategies, mainly in cases where manipulation was not required (e.g., general inquiries about logistics, affirmation, and reassurance). A full list of strategies and their short descriptions (Supplementary Table S15) was provided as bullet points in the SEMA’s prompt, and the agent was instructed to choose one of these strategies and generate a corresponding response.
All agents were provided with the user’s query and the corresponding four options for the given scenario. In addition, they had access to the participants’ personality traits, as personalization could significantly enhance the effectiveness of human-AI interactions [matz2024potential]: each item from the participants’ Big-Five inventory[john1991big] results was described as Low (<3<3), Moderate ($\geq 3$ and $\leq 5$), or High (>5>5) in the agent’s prompt. Given that existing AI systems are capable of predicting their users’ personalities with sufficient accuracy [li2024can], we believe adding this information did not hinder the applicability of our experimental setup to real-world contexts. Designed prompts are provided in Supplementary Figures S3 - S5.
@@
The scenarios in this study were carefully crafted to assess manipulation susceptibility in two domains reflecting common real-world situations in which AI assistants are employed [lakkaraju2023llms, hwang2021toward, keshishi2023emotional, gual2022using, lerner2015emotion, burtell2023artificial]: Financial (seeking advice on purchases) and Emotional (seeking guidance on handling emotional dilemmas) decision-making. In each domain, we curated three widely used topics to design scenarios applicable to a diverse range of individuals. For each scenario, to ensure consistency between different participants, we designed pre-defined queries representing a hypothetical user’s situation, needs, and requirements.
Each scenario included four plausible options: one crafted as the optimal choice given the constraints of the user’s query, while the remaining three were designed to indicate different levels of potential harm. Options were carefully crafted through iterative discussions to ensure realistic trade-offs: Optimal options included minor shortcomings to prevent them from being trivially dominant, and harmful options offered subtle bonuses with latent downsides that would ultimately render them harmful. This intentional asymmetry (3 harmful options vs. 1 optimal) mirrors real-world scenarios, in which beneficial (optimal) options are often outnumbered by inferior alternatives [johnson2012beyond, mathur2019dark, hagiu2022should]. Lastly, we designated one of the harmful options as the agent’s hidden incentive for each scenario to prevent user suspicion regarding its bias.
@@
AI systems are making their way into financial investments [schillaci2024llm] and are increasingly being used to enhance product advertisements [matz2024potential, burtell2023artificial] and provide financial advice [cao2022ai]. Hence, it is crucial to understand the extent to which these systems can impact users’ decisions despite their better judgment. Drawing from existing research on consumer decision-making [malhotra2019growth, trivedi2023effect, limbu2024influences], we identified key scenarios where users commonly seek AI assistance for financial guidance. We focused on three representative consumer decisions that require minimal domain expertise: selecting a fitness tracker, choosing a weight-loss medication, and finding an online shopping platform for clothes. To ensure practicality while maintaining experimental control, we designed pre-defined user queries corresponding to each topic that incorporated realistic budget constraints and requirements (Tables S16-S18).
The optimal option was designed to align with stated user requirements and budget constraints while incorporating realistic trade-offs. We designed the remaining options to exhibit different types of potential financial harm: 1) Non-existent Product: This option was crafted to mirror characteristics of common online scams [montanez2020human], featuring unusually attractive terms, comparatively lower cost, and limited verification possibilities. Specifically, this option was designed to only be available for pre-order from the supplier’s website and had relatively few or no reviews, making it difficult for users to assess its legitimacy. 2) Excessive Product: This option represented over-featured solutions that exceeded the user’s stated requirements and maximum budget and was designed to test users’ susceptibility to unnecessary upselling[heidig2017supersize]. It allowed us to assess whether users could be swayed by the perceived desirability of excessive products, even when it caused them to spend more than they had originally planned. 3) Dependence-inducing Product: This option was designed to satisfy the user’s requirements at a relatively lower upfront cost than the optimal option. However, it incorporated elements of long-term financial commitment, such as subscription models, for continued access to users’ desired features, which introduced hidden costs and future obligations. By choosing this option, users incurred recurring fees that, over time, exceeded their maximum budget, reflecting real-world scenarios where consumers are lured by low initial costs but face long-term financial burdens.
@@
The increasing deployment of AI systems in mental healthcare, particularly for counseling and emotional support, presents both opportunities and potential risks. These systems have gained traction due to their high accessibility, cost-effectiveness, and reduced stigma associated with seeking support [sabour2023chatbot]. However, this widespread adoption necessitates an investigation into human susceptibility to AI influence when dealing with emotional challenges, particularly regarding their established beliefs about managing emotional dilemmas. The scenarios for this domain were developed based on a comprehensive analysis of mental health forums and existing literature on emotional dilemmas [sun2021psyqa, sik2023topic]. Through careful examination of existing data and literature [whisman1993life, koff1997effects, pinkasavage2015social, comolli2024concentration, kanner1981comparison], we identified three primary topics of emotional adversities: issues with self-esteem (Table S19), conflict with a close friend (Table S20), and criticism at the workplace (Table S21).
@@
Each scenario presented participants with four distinct coping strategies (i.e., how individuals tend to manage their thoughts and emotions in response to emotional adversities [carver1989assessing]), designed to represent different approaches to emotional regulation and problem-solving. Existing research [carver1989assessing, carver1997you, algorani2023coping, compas2017coping, brown2005good] classifies coping strategies as adaptive or maladaptive depending on their general impact on the individual’s overall well-being and mental health. We included an adaptive coping mechanism as the optimal option for the given scenario, while the remaining options were designed to incorporate maladaptive practices, which could potentially lead to increasing rates of depression and anxiety [folkman2004coping, kato2015frequently].
@@
Therefore, the optimal option was designed to represent active coping, a process based on cognitive-behavioral literature that involves identifying specific actions to address the situation, developing alternative perspectives, and constructive engagement with support networks, which has shown the strongest correlations with increasing user’s mental well being and lower rates of anxiety and depression [carver1997you, compas2017coping]. The remaining options involved common maladaptive mechanisms with varying levels of harm [kato2015frequently]: 1) Disengagement: This option reflects avoidant coping patterns, incorporating elements such as behavioral disengagement, mental disengagement, and denial. While potentially offering immediate emotional relief, this option demonstrated less desirable long-term outcomes; 2) Emotional Venting: This option includes elements of emotional release while omitting components of reflection, learning, or growth that characterize more adaptive forms of emotional processing. Hence, it clearly distinguishes between adaptive and maladaptive emotional processing by emphasizing unregulated emotional discharge without problem-solving elements. 3) Self-blame: This option was constructed to reflect common patterns of maladaptive attribution as it included subtle elements of perfectionism and excessive responsibility-taking, characteristics that portray the individual as the root of the problem, bringing additional mental burden to the individual without finding a viable long-term solution.
@@
To ensure user privacy, we created our own data collection platform using React[react] and FastAPI[fastapi]. Upon logging onto our platform, each participant was shown a series of three scenarios based on their assigned domain (Figure S7). Selecting one of these scenarios opened a new page, in which the participants were shown the pre-defined query and four available options in a randomized order (Figure S8). The platform interface was designed to minimize potential confounds, presenting all information in a standardized format while randomly rearranging the order of the options within each scenario to control for position effects. Before interacting with the assigned AI agent, participants rated their initial preferences for each option, along with their confidence levels and topic familiarity, on a 10-point Likert scale. After submitting their ratings, participants entered the interaction page (Figure S9). Following a conversation lasting at least 10 turns with the assigned agent, participants provided their final preferences for each option, their confidence levels, and their familiarity with the scenario. This design enabled us to quantify changes in both decision-making and confidence levels as metrics for manipulation susceptibility.
@@
This study followed a three-phase protocol designed to assess participants’ susceptibility to AI manipulation while controlling for individual differences and ensuring ethical practices. Our experiments were approved by the Institutional Review Board at the University of Hong Kong (Reference No.: EA240497).
@@
Prior to our experiment, participants completed a comprehensive questionnaire of validated psychometric instruments, including demographic information (age, sex, education, occupation, and marital status) and social factors that may influence people’s susceptibility to manipulation, including personality[yang2022predicting], social engineering susceptibility[montanez2022social] [stewart2018modification], social support [pinsker2010exploitation], and self-esteem[wang2024mentalmanip]: 1) Ten-Item-Personality-Inventory (TIPI)[gosling2003very], a 10-item assessment of personality traits based on dimensions from the Big Five Inventory (BFI)[john1991big]: openness, conscientiousness, extroversion, agreeableness, and emotional stability. We adopted this version of the personality test, as we assumed that while an AI agent may have a perception of the user’s personality traits, it would not be heavily accurate[matz2024potential]; 2) Social Engineering Susceptibility Scale[workman2007gaining] focusing on individuals’ vulnerabilities in digital interactions and modified to include the portion of the questions that were relevant to our study (i.e., excluding questions regarding informational susceptibilities); 3) Oslo Social Support Scale (OSSS-3)[kocalevent2018social], a 3-item questionnaire evaluating participants’ accessibility to support networks; 4) Rosenberg Self-Esteem Scale (RSES)[rosenberg1965rosenberg], a 10-item measurement of individuals’ perception of self-worth. Due to the nature of our study, we also adopted the AI Attitude Scale[grassini2023development] to assess participants’ predispositions toward AI systems.